(b) [5 marks] Using first principles, derive the radioactive decay rate for nuclei whose decay constant is ?. Here R,-??, is the rate of decay at t-0 [5 marks] A count-rate meter is used to measure the activity of a given sample. At one instance it shows 4750 counts per minute (cpm). Five minutes later it shows 2700 cpm. Find the decay constant and the half-life of the sample. (c) (d) [5 Marks] For alpha decay, show that 13 markel Thus for 212 Ri ?208 TI He+O calculate K
Answer #1

Lt initi all R, ณ.tn.be lke กัmbun, of.Tud:0 a ctive elamwe Vn decor eleme Кос เกี่ and timed num ber decrens 껀 cons-^avi iniKI = 4750 60入 0) -L50R. e -300 λ 一60 475φ 2구。 300 λ 240 17593 e 240 λ = log( 1.7599) =0.5649 λ - 0.5649 2t D = 2.35xt03-0,002conserv Khon rt energi dV4b, ves ne kinetc Puotiele212 ー→ 20 8 3 81 Am- 212-o 98040-208.088-4.0026 = 0.00744 62GG4 Mev A·00 26 =0,019235 20g o 88 2664 MRN ol 235) RN o1923 S